# Semantic Scene Completion Based 3D Traversability Estimation for Off-Road Terrains

### v1

@@ V Conclusion
In this paper, We propose ORDformer, a multimodal method for off-road traversability estimation that utilizes LiDAR point clouds and monocular images to generate dense semantic occupancy predictions. By integrating geometric and semantic information, ORDformer enhances environmental feature extraction, enabling more accurate estimation of traversable regions in complex, occluded terrains. Furthermore, we introduce the RELLIS-OCC dataset with 3D traversable occupancy annotations, enabling detailed terrain assessment in off-road environments. Extensive experiments validate the superior performance of ORDformer compared to existing approaches, particularly in handling obstacles with diverse geometries and partial occlusion. ORDformer holds significant potential for real-world deployment, as it can be easily adapted to different vehicle platforms and a range of off-road environments, providing a robust solution for autonomous navigation in challenging conditions.
For future work, the proposed traversability estimation method will be integrated with 3D planning algorithms to enable comprehensive 3D trajectory planning in off-road conditions. Additionally, recognizing the limitations of existing public datasets, we plan to collect an expanded off-road dataset containing complex obstacle information and evaluate the performance of ORDformer in detecting negative and overhanging obstacles.

### v2

@@ V Conclusion
In this paper, we propose 3DTTNet, a multimodal method for off-road traversable terrain modeling that utilizes LiDAR point clouds and monocular images to generate 3D traversable predictions. On basis of combining geometric and semantic information, 3DTTNet enhances environmental feature extraction, enabling more accurate estimation of traversable regions in complex, occluded terrains. Furthermore, we introduce the RELLIS-OCC dataset with 3D traversable annotations, enabling detailed terrain assessment in off-road environments. Extensive experiments validate the superior performance of 3DTTNet compared to baseline approaches, particularly in handling obstacles with diverse geometries and partial occlusion. 3DTTNet holds significant potential for real-world deployment, as it can be easily adapted to different platforms and a range of off-road environments, providing a robust solution for autonomous navigation in challenging conditions.
For future work, the proposed traversability estimation framework will be united with 3D motion planning algorithms to enable autonomous navigation in off-road conditions. Additionally, recognizing the limitations of existing public datasets, we plan to collect an expanded off-road dataset containing complex obstacle information and evaluate the performance of 3DTTNet in detecting negative and overhanging obstacles.